# Quantum variational calculus on a lattice

### v1

@@ Abstract.
We solve the long-standing problem of variational calculus on a noncommutative space or spacetime for a significant class of models with trivial jet bundle. Our approach entails a quantum version of the Anderson variational double complex $\Omega(J^{\infty})$ and includes Euler-Lagrange equations and a partial Noether’s theorem. We show in detail how this works for a free field on a $\mathbb{Z}^{m}$ lattice regarded as a discrete noncommutative geometry, obtaining the Klein-Gordon equation for a scalar field, including with a general metric and gauge field background, as the Euler-Lagrange equations of motion for an action. In the case of a flat metric we also obtain an exactly on-shell conserved stress-energy tensor and Noether charges for a scalar field on the lattice.
@@ 1. Introduction
What has been missing for over 30 years, however, is the single most important thing needed for a convincing derivation of physics on a noncommutative space or spacetime, including on a lattice as an exact geometry. Namely, a formulation of variational calculus whereby an action functional as used in a path integral formulation of the quantum field theory on such spaces is connected to equations of motion on such spaces. Without this, current attempts to formulate physics, while promising, lack a complete picture even at the level of classical field theory let alone quantum field theory. In the present work, we propose a solution to this long-standing problem for a limited class of noncommutative geometries but one that includes the lattice case.
@@
$E[\phi]=-{1\over 2}(\partial_{+}\phi)(\partial_{-}\phi)+{m^{2}\over 2}\phi^{2}$
of a free scalar field of mass $m$ on a lattice line $\mathbb{Z}$ appears to be new (this is the simplest case of Corollary 4.3). Here, (±$(\partial_{\pm}\phi)(i)=\phi(i\pm 1)-\phi(i)$ and the claim is that $E[\phi]$ is constant on $\mathbb{Z}$ if $\phi$ obeys the discrete wave equation $(\Delta_{\mathbb{Z}}+m^{2})\phi=0$, where $\Delta_{\mathbb{Z}}\phi=\partial_{+}+\partial_{-}$ is the usual discrete Laplacian. On a literature search, we noted [6], which does not appear to have an exactly conserved stress tensor.
An outline of the paper is as follows. Since the paper involves a number of less familiar methods, including the variational double complex, we will build up the theory in several layers starting with a preliminary Section 2 explaining in detail how everything proceeds in the case of classical field theory on a classical manifold $M$. Section 3 then gives the discrete version where $M$ is replaced by a lattice line $\mathbb{Z}$, as the easiest case of our theory. Section 4 then covers $\mathbb{Z}^{m}$ and $\mathbb{Z}^{1,m-1}$ (i.e. Euclidean and Minkowski lattice cases) and more generally for the ‘spacetime’ any discrete group $G$ provided we can construct the jet bundle (this includes all Abelian groups but in principle also some nonAbelian groups such as the group $S_{3}$ of permutations of three elements [20]). Section 5 looks at scalar fields with a general metric and/or $U(1)$ gauge field on the lattice as background, using methods of quantum Riemannian geometry[2] and lattice gauge theory[21]. We end with some concluding remarks in Section 6 about further work.
@@ 2.1. Jet bundle and variational bicomplex
We focus on $E=M\times\mathbb{R}\to M$ as the bundle for matter fields so that $\Gamma(E)=C^{\infty}(M)=F$ is the ‘space of matter fields’ as a linear space, where a function $f\in F$ is viewed as a map $M\to E$ sending $x\mapsto(x,\phi(x))$. The jet bundle here is trivial and this allows us to write jet prolongation map $j_{\infty}\colon F\to\Gamma(J^{\infty})$ explicitly as
@@
$e_{\infty}^{*}(\mathrm{d}_{H}\Phi)={\partial_{i}}\big{(}\Phi(j_{\infty}(\cdot)(\cdot))\big{)}\mathrm{d}x^{i}=\left(\left({\partial_{i}\Phi}\right)(j_{\infty}(\cdot)(\cdot))+\left({\partial\Phi\over\partial u_{I}}\right)(j_{\infty}(\cdot)(\cdot))\partial_{iI}(\cdot)(\cdot)\right)\mathrm{d}x^{i}$
where we evaluated at $f\in F$ and use the chain rule. However, we can compute
@@ 2.2. Euler-Lagrange equations
Computing the RHS gives $\mathrm{d}_{V}L\wedge\mathrm{Vol}=EL-\mathrm{d}_{h}\Theta$ for $EL\in\Omega^{{\mathrm{top}},1}(J^{\infty})$ the EL form and $\Theta\in\Omega^{{\mathrm{top}}-1,1}(J^{\infty})$ the boundary term. Explicitly, we have
@@ 2.3. Symmetries and Noether’s Theorem
$X_{\infty}=X^{i}\partial_{i}+X^{a}_{I}\frac{\partial}{\partial u^{a}_{I}},\qquad X_{I}^{a}=D_{I}(X^{a}-X^{i}u^{a}_{i})+X^{i}u^{a}_{iI}.$
This can be split into horizontal and vertical components $X_{\infty}=X_{H}+X_{V}$ with respect to the contact structure,
@@
such that $[\iota_{X_{H}},\mathrm{d}_{V}]=[\iota_{X_{V}},\mathrm{d}_{H}]=0$ and
@@
Following [8, Chap. 2.6], one says that $X_{E}$ is a symmetry if there is a form $\sigma_{X}\in\Omega^{{\mathrm{top}}-1,0}(J^{\infty})$ such that
@@
cf. [8, Chap. 2.6] Given a symmetry $(X_{E},\sigma_{X})$, there is an associated (on-shell) conserved current $j_{X}:=\sigma_{X}-\iota_{X_{H}}(L\mathrm{Vol})-\iota_{X_{V}}\Theta\in\Omega^{{\mathrm{top}}-1,0}(J^{\infty})$. Then
@@
where we have used $[\iota_{X_{H}},\mathrm{d}_{V}]=[\iota_{X_{V}},\mathrm{d}_{H}]=0$, $\mathcal{L}_{X}=\mathcal{L}_{X_{H}}+\mathcal{L}_{X_{V}}$ and $\mathcal{L}_{X_{V}}(L\mathrm{Vol})=\iota_{X_{V}}\mathrm{d}_{V}(L\mathrm{Vol})=\iota_{X_{V}}EL-\iota_{X_{V}}\mathrm{d}_{H}\Theta$. ∎
@@
The quantity $Q=\int_{\Sigma}e_{\infty}^{*}(\cdot,\phi)(j^{0}_{X})\mathrm{Vol}_{\Sigma}$ is conserved in time in the sense that $\partial_{0}Q$ vanishes on-shell. $Q$ is called a conserved charge.
@@
$\partial_{0}Q=-\int_{\Sigma}e_{\infty}^{*}(\cdot,\phi)(D_{i}j^{i}_{X})\mathrm{Vol}_{\Sigma}=-\int_{\Sigma}\partial_{i}(e_{\infty}^{*}(\cdot,\phi)j^{i}_{X})\mathrm{Vol}_{\Sigma}=0$
as $\Sigma$ does not have a boundary. ∎
@@
$E[\phi]={1\over 2}\int_{\Sigma}\left((\partial_{0}\phi)^{2}-g_{ii}\sum^{\dim\Sigma}_{i=1}(\partial_{i}\phi)^{2}+m^{2}\phi^{2}\right)\mathrm{d}^{n-1}x,$
@@
as expected. As in Example 2.6, this Lagrangian is translation invariant, and a similar analysis results in
$T=(\bar{u}_{i}u_{j}-g_{ij}L)\mathrm{d}x^{i}\otimes_{C^{\infty}(M)}\mathrm{d}x^{j}.$
In terms of the field $\phi$ the conserved $U(1)$-current and the stress-energy tensor read
$j_{U(1)}[\phi]=-\frac{{\mathfrak{i}}\varphi}{2}g^{im}((\partial_{m}\phi^{*})\phi-\phi^{*}(\partial_{m}\phi))\mathrm{Vol}_{i},$
$T[\phi]=(\partial_{i}\phi^{*}\partial_{j}\phi-g_{ij}L)\mathrm{d}x^{i}\otimes_{C^{\infty}(M)}\mathrm{d}x^{j}.$
The conserved charge corresponding to the $U(1)$ symmetry in the case of $M=\mathbb{R}\times\Sigma$ with Euclidean or Minkowski metrics is then
$Q_{U(1)}=-\frac{{\mathfrak{i}}\varphi}{2}g^{00}\int_{\Sigma}((\partial_{0}\phi^{*})\phi-\phi^{*}(\partial_{0}\phi))\mathrm{d}^{n-1}x.$
@@ 3. Variational calculus on the integer lattice $\mathbb{Z}$
We now apply the exact same ideology but replacing the base manifold $M$ with $\mathbb{Z}$ as a discrete noncommutative geometry. There is a more general picture whereby $\Omega^{1}$ on a discrete set is precisely a graph, and here we consider $\mathbb{Z}$ as a graph, namely a one-dimensional lattice. Using that $\mathbb{Z}$ is a group, it is convenient to write the calculus in terms of a basis of invariant 1-forms $e^{a}=e^{\pm}$ over the algebra of real valued functions $C(\mathbb{Z})$. The exterior derivative is then
@@
±$\partial_{\pm}(fg)=(\partial_{\pm}f)R_{\pm}(g)+f\partial_{\pm}(g)$
for all $f,g\in C(Z)$. Higher forms have a basis given by that of the Grassmann algebra on the $e^{a}$ (they anticommute among themselves) and $\mathrm{d}=[\theta,\ \}$ where $\theta=\sum_{a}e^{a}$ and [,}[\ ,\ \}[ , } denotes a graded commutator. This is how lattice geometry appears as an example of noncommutative geometry. It has been used in numerous works, see [5, 18, 4, 2], as an example of a general analysis for differentials on quantum groups [25]. In our case, there is also a flat torsion-free connection $\nabla\colon\Omega^{1}\to\Omega^{1}\otimes_{C(\mathbb{Z})}\Omega^{1}$ characterised by $\nabla e^{\pm}=0$ (it is the QLC for the constant metric in the lattice in the sense of [2].) Notice that $\Omega^{1}(\mathbb{Z})$ is 2-dimensional over the algebra because $\partial_{\pm}$ are linearly independent as operators. For the same reason the top ‘volume’ form ${\rm Vol}:=e^{+}\wedge e^{-}$ is a 2-form. We fix the above calculus $\Omega(\mathbb{Z})$ and our first task is to extend it to a calculus $\Omega(J^{\infty})$.

### v2

@@ Abstract.
We solve the long-standing problem of variational calculus on a noncommutative space or spacetime for a significant class of models with trivial jet bundle. Our approach entails a quantum version of the Anderson variational double complex $\Omega(J^{\infty})$ and includes Euler-Lagrange equations and a partial Noether’s theorem. We show in detail how this works for a free field on a $\mathbb{Z}^{m}$ lattice regarded as a discrete noncommutative geometry, obtaining the Klein-Gordon equation for a scalar field, including with a general metric and gauge field background, as the Euler-Lagrange equations of motion for an action. In the case of a flat metric we also obtain an exactly on-shell conserved stress-energy tensor and Noether charges for a scalar field on the lattice and modified energy-momentum relations.
@@ 1. Introduction
What has been missing for over 30 years, however, is the single most important thing needed for a convincing derivation of physics on a noncommutative space or spacetime, including on a lattice as an exact geometry. Namely, a formulation of variational calculus whereby an action functional as used in a path integral formulation of the quantum field theory on such spaces is connected to equations of motion on such spaces and to conserved quantities. Without this, current attempts to formulate physics, while promising, lack a complete picture even at the level of classical field theory let alone quantum field theory. In the present work, we propose a solution to this long-standing problem for a limited class of noncommutative geometries but one that includes the lattice case.
@@
(1) $E[\phi]=-{1\over 2}(\partial_{+}\phi)(\partial_{-}\phi)+{m^{2}\over 2}\phi^{2}$
of a free scalar field of mass $m$ on a lattice line $\mathbb{Z}$ appears to be new (this is the simplest case of Corollary 4.3). Here, $(\partial_{\pm}\phi)(i)=\phi(i\pm 1)-\phi(i)$ and the claim is that $E[\phi]$ is constant on $\mathbb{Z}$ if $\phi$ obeys the discrete wave equation $(\Delta_{\mathbb{Z}}+m^{2})\phi=0$, where $\Delta_{\mathbb{Z}}\phi=\partial_{+}+\partial_{-}$ is the usual discrete Laplacian. There are bounded oscillatory solutions here for all real values of $E[\phi]$ and $m<2$ relative to the discretization scale but real solutions only for $E[\phi]>0$. We will look at this look this from the point of view of a discrete time harmonic oscillator as a classical particle $q\colon\mathbb{Z}\to\mathbb{R}$ with frequency $\omega$ playing the role of $m$, see Figure 1, and from a scalar field theory, see Figures 2-5. On a literature search, we noted [6], but this does not appear to have an exactly conserved stress tensor.
An outline of the paper is as follows. Since the paper involves a number of less familiar methods, including the variational double complex, we will build up the theory in several layers starting with a preliminary Section 2 explaining in detail how everything proceeds in the case of classical field theory on a classical manifold $M$. Section 3 then gives the discrete version where $M$ is replaced by a lattice line $\mathbb{Z}$, as the easiest case of our theory. Section 4 then covers $\mathbb{Z}^{m}$ and $\mathbb{Z}^{1,m-1}$ (i.e. Euclidean and Minkowski lattice cases) and more generally for the ‘spacetime’ any discrete group $G$ provided we can construct the jet bundle. This includes all Abelian groups but in principle also some nonAbelian groups such as the group $S_{3}$ of permutations of three elements [21]. Section 4.2 focusses on (1+1)-dimensional lattice scalar fields in both the Euclidean and Minkowski cases, showing modified energy-momentum relations. For example, in the Minkowski case we have
$\mathcal{E}^{2}+\left(\frac{m^{2}}{2}-\sqrt{1-\mathcal{P}^{2}}\right)^{2}-1=0$
for the solution that is close to the continuum case. Note that even in the continuum [8], classical variational calculus does not lead to the familiar form in quantum field theory as this requires quantisation, but one can get to the familiar form assuming an energy density per quantum. We used a similar strategy in the lattice case pending a treatment of quantum field theory on the lattice. Section 5 looks at scalar fields with a general metric and/or $U(1)$ gauge field on the lattice as background, using methods of quantum Riemannian geometry [2] and lattice gauge theory [22]. We end with some concluding remarks in Section 6 about further work.
@@ 2.1. Jet bundle and variational bicomplex
We focus on $E=M\times\mathbb{R}\to M$ as the bundle for matter fields so that $\Gamma(E)=C^{\infty}(M)=F$ is the ‘space of matter fields’ as a linear space, where a function $\phi\in F$ is viewed as a map $M\to E$ sending $x\mapsto(x,\phi(x))$. The jet bundle here is trivial and this allows us to write the jet prolongation map $j_{\infty}\colon F\to\Gamma(J^{\infty})$ explicitly as
@@
$e_{\infty}^{*}(\mathrm{d}_{H}\Phi)={\partial_{i}}\big(\Phi(j_{\infty}(\cdot)(\cdot))\big)\mathrm{d}x^{i}=\left(\left({\partial_{i}\Phi}\right)(j_{\infty}(\cdot)(\cdot))+\left({\partial\Phi\over\partial u_{I}}\right)(j_{\infty}(\cdot)(\cdot))\partial_{iI}(\cdot)(\cdot)\right)\mathrm{d}x^{i},$
where we evaluated at $\phi\in F$ and use the chain rule. However, we can compute
@@ 2.2. Euler-Lagrange equations
Computing the RHS gives $\mathrm{d}_{V}L\wedge\mathrm{Vol}=EL-\mathrm{d}_{H}\Theta$ for certain $EL\in\Omega^{{\mathrm{top}},1}(J^{\infty})$ the EL form and $\Theta\in\Omega^{{\mathrm{top}}-1,1}(J^{\infty})$ the boundary term. Explicitly, we have
@@ 2.3. Symmetries and Noether’s Theorem
$X_{\infty}=X^{i}\partial_{i}+X^{a}_{I}\frac{\partial}{\partial u^{a}_{I}},\qquad X_{I}^{a}=D_{I}(X^{a}-X^{i}u^{a}_{i})+X^{i}u^{a}_{iI},$
where $D_{I}=D_{i_{1}}\cdots D_{i_{k}}$ for the multiindex $I=\{i_{1},\cdots,i_{k}\}$. This can be split into horizontal and vertical components $X_{\infty}=X_{H}+X_{V}$ with respect to the contact structure,
@@
such that $\iota_{X_{H}},\mathrm{d}_{V}]=[\iota_{X_{V}},\mathrm{d}_{H}]=0$ and
@@
Following [8, p.165ff], one says that $X_{E}$ is a symmetry if there is a form $\sigma_{X}\in\Omega^{{\mathrm{top}}-1,0}(J^{\infty})$ such that
@@
cf. [8, p.165ff] Given a symmetry $(X_{E},\sigma_{X})$, there is an associated (on-shell) conserved current $j_{X}:=\sigma_{X}-\iota_{X_{H}}(L\mathrm{Vol})-\iota_{X_{V}}\Theta\in\Omega^{{\mathrm{top}}-1,0}(J^{\infty})$. Then
@@
where we have used $\iota_{X_{H}},\mathrm{d}_{V}]=[\iota_{X_{V}},\mathrm{d}_{H}]=0$, $\mathcal{L}_{X}=\mathcal{L}_{X_{H}}+\mathcal{L}_{X_{V}}$ and $\mathcal{L}_{X_{V}}(L\mathrm{Vol})=\iota_{X_{V}}\mathrm{d}_{V}(L\mathrm{Vol})=\iota_{X_{V}}EL-\iota_{X_{V}}\mathrm{d}_{H}\Theta$. ∎
@@
The quantity $Q[\phi]=\int_{\Sigma}e_{\infty}^{*}(j^{0}_{X})(\cdot,\phi)\mathrm{Vol}_{\Sigma}$ is conserved in time in the sense that $\partial_{0}Q$ vanishes on-shell. $Q$ is called a conserved charge. The quantities $\rho[\phi]\coloneqq e_{\infty}^{*}(j^{0}_{X})(\cdot,\phi)$ and $J^{i}[\phi]\coloneqq e_{\infty}^{*}(j^{i}_{X})(\cdot,\phi)$ are called charge and current densities, and satisfy the continuity equation $\partial_{0}\rho+\partial_{i}J^{i}=0$ when $\phi$ is on-shell.
@@
$\partial_{0}Q[\phi]=-\int_{\Sigma}e_{\infty}^{*}(D_{i}j^{i}_{X})(\cdot,\phi)\mathrm{Vol}_{\Sigma}=-\int_{\Sigma}\partial_{i}(e_{\infty}^{*}(j^{i}_{X})(\cdot,\phi))\mathrm{Vol}_{\Sigma}=0$
as $\Sigma$ does not have a boundary. The continuity equation can be derived by directly applying $e^{*}_{\infty}$ to $\mathrm{d}_{H}j_{X}$. ∎
@@
$E[\phi]=\int_{\Sigma}{1\over 2}\left((\partial_{0}\phi)^{2}-g_{ii}\sum^{\dim\Sigma}_{i=1}(\partial_{i}\phi)^{2}+m^{2}\phi^{2}\right)\mathrm{d}^{n-1}x,$
@@
Focusing now on the (1+1)-case for simplicity and choosing a plane wave solution $\phi(t,x)=A\cos(\omega t+g_{11}\kappa x)$ solving the Klein-Gordon equation $(\Delta+m^{2})\phi=0$ with the dispersion relations $m^{2}=\omega^{2}+g_{11}\kappa^{2}$, one recovers that the expected energy and momentum densities (the integrands of $E[\phi]$ and $P_{x}[\phi]$) are given by
$\mathcal{E}_{dens}[\phi]={A^{2}\over 2}\left(\omega^{2}\sin^{2}(\omega t+g_{11}\kappa x)-g_{11}\kappa^{2}\sin^{2}(\omega t+g_{11}\kappa x)+m^{2}\cos^{2}(\omega t+g_{11}\kappa x)\right),$
$\mathcal{P}_{x,dens}[\phi]=A^{2}g_{11}\omega\kappa\sin^{2}(\omega t+g_{11}\kappa\cdot x).$
As is common in real scalar field theory, it is not directly clear that integrating these densities over $\Sigma$ will result in a time independent result. Instead, we can look at their spatial average over $x_{0},x_{0}+2\pi/\kappa]$ to find [8, p.194f]
$\langle\mathcal{E}_{dens}[\phi]\rangle=\frac{\kappa}{2\pi}\int^{x_{0}+2\pi/\kappa}_{x_{0}}\mathcal{E}_{dens}\,\mathrm{d}x={A^{2}\over 2}\omega^{2},$
$\langle\mathcal{P}_{x,dens}[\phi]\rangle=\frac{\kappa}{2\pi}\int^{x_{0}+2\pi/\kappa}_{x_{0}}\mathcal{P}_{x,dens}\,\mathrm{d}x=g_{11}{A^{2}\over 2}\omega\kappa,$
which makes the time independence of the energy and momenta clear. Using the dispersion relation we find that these densities are related as
$\langle\mathcal{E}_{dens}[\phi]\rangle^{2}+g_{11}\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}={A^{2}\over 2}\langle\mathcal{E}_{dens}[\phi]\rangle\,m^{2},$
and choosing the normalization $A=\sqrt{2/\langle\mathcal{E}_{dens}[\phi]\rangle}$ of the field $\phi$ we then recover the familiar $\langle\mathcal{E}_{dens}[\phi]\rangle^{2}+g_{11}\langle\mathcal{P}_{x,dens}[\phi]\rangle^{2}=m^{2}$, indicating that the field $\phi$ has the density of ‘one particle of mass $m$ per unit volume’, with $\langle\mathcal{E}_{dens}[\phi]\rangle=\omega$ and $\langle\mathcal{P}_{x,dens}[\phi]\rangle=\kappa$. Note that the usual normalization $\sqrt{1/(2\langle\mathcal{E}_{dens}[\phi]\rangle)}$ found when quantizing the field via the CCR is recovered if we write the field $\phi(t,x)=(A/2)(e^{{\mathfrak{i}}(\omega t+g_{11}\kappa x)}+e^{-{\mathfrak{i}}(\omega t+g_{11}\kappa x)})$ in terms of exponentials as it is done when quantizing it.
@@
as expected. In terms of the field $\phi$ and its conjugate the conserved $U(1)$-current reads
$j_{U(1)}[\phi,\phi^{*}]=-\frac{{\mathfrak{i}}\varphi}{2}g^{im}((\partial_{m}\phi^{*})\phi-\phi^{*}(\partial_{m}\phi))\mathrm{Vol}_{i}.$
and for $M=\mathbb{R}\times\Sigma$, we find the conserved charge corresponding to the $U(1)$-symmetry to be
$Q_{U(1)}[\phi,\phi^{*}]=-\frac{{\mathfrak{i}}\varphi}{2}g^{0m}\int_{\Sigma}((\partial_{m}\phi^{*})\phi-\phi^{*}(\partial_{m}\phi))\mathrm{d}^{n-1}x.$
with charge density
$J^{m}_{U(1)}[\phi,\phi^{*}]=-\frac{{\mathfrak{i}}\varphi}{2}g^{mk}((\partial_{k}\phi^{*})\phi-\phi^{*}(\partial_{k}\phi)).$
As in Example 2.6, this Lagrangian is translation invariant for $g$ the Euclidean or Minkowski metrics, and a similar analysis results in the expressions
$T=(\bar{u}_{i}u_{j}-g_{ij}L)\mathrm{d}x^{i}\otimes_{C^{\infty}(M)}\mathrm{d}x^{j},$
$T[\phi,\phi^{*}]=(\partial_{i}\phi^{*}\partial_{j}\phi-g_{ij}L)\mathrm{d}x^{i}\otimes_{C^{\infty}(M)}\mathrm{d}x^{j}.$
We again focus on the (1+1)-dimensional case, where a plane wave solution $\phi(t,x)=Ae^{-{\mathfrak{i}}(\omega t+g_{11}\kappa x)}$ of the Klein-Gordon equation satisfies the dispersion relation $m^{2}=\omega^{2}+g_{11}\kappa^{2}$. Then the energy, momentum, $U(1)$-charge and $U(1)$-current densities are recovered as expected by [8, p.194f]
$\mathcal{E}_{dens}[\phi,\phi^{*}]=T_{00}[\phi,\phi^{*}]=|A|^{2}\omega^{2},$
$\mathcal{P}_{x,dens}[\phi,\phi^{*}]=T_{01}[\phi,\phi^{*}]=g_{11}|A|^{2}\omega\kappa,$
$\rho_{U(1),dens}[\phi,\phi^{*}]=|A|^{2}\,\varphi\omega,$
$J_{U(1),dens}[\phi,\phi^{*}]=|A|^{2}\,\varphi\kappa.$
Note that $\mathcal{E}_{dens}[\phi,\phi^{*}]$, $\mathcal{P}_{x,dens}[\phi,\phi^{*}]$ are double the values of the real scalar field theory in Example 2.6 due to the existence of two different modes $\phi$, $\phi^{*}$ in this theory. In this case we find the relations
$\mathcal{E}_{dens}[\phi,\phi^{*}]^{2}+g_{11}\mathcal{P}_{x,dens}[\phi,\phi^{*}]^{2}=|A|^{2}\mathcal{E}_{dens}[\phi,\phi^{*}]\,m^{2},$
directly for the energy and momentum densities, which with $A=\sqrt{1/\mathcal{E}_{dens}[\phi,\phi^{*}]}$ becomes
$\mathcal{E}_{dens}[\phi,\phi^{*}]^{2}+g_{11}\mathcal{P}_{x,dens}[\phi,\phi^{*}]=m^{2}.$
@@ 3. Variational calculus on the integer lattice $\mathbb{Z}$
We now follow the exact same ideology but replacing the base manifold $M$ with $\mathbb{Z}$ as a discrete noncommutative geometry. This is an application of a general picture whereby $\Omega^{1}$ on a discrete set is precisely a graph, and here we consider $\mathbb{Z}$ as a graph, namely a one-dimensional lattice. Using that $\mathbb{Z}$ is a group, it is convenient to write the calculus in terms of a basis of invariant 1-forms $e^{a}=e^{\pm}$ over the algebra of real valued functions $C(\mathbb{Z})$. The exterior derivative is then
@@
$\partial_{\pm}(fg)=(\partial_{\pm}f)R_{\pm}(g)+f\partial_{\pm}(g)$
for all $f,g\in C(Z)$. Higher forms have a basis given by that of the Grassmann algebra on the $e^{a}$ (they anticommute among themselves) and $\mathrm{d}=[\theta,\ \}$ where $\theta=\sum_{a}e^{a}$ and [,}[\ ,\ \} denotes a graded commutator. This is how lattice geometry appears as an example of noncommutative geometry. It has been used in numerous works, see [5, 19, 4, 2], as an example of a general analysis for differentials on quantum groups [28]. In our case, there is also a flat torsion-free connection $\nabla\colon\Omega^{1}\to\Omega^{1}\otimes_{C(\mathbb{Z})}\Omega^{1}$ characterised by $\nabla e^{\pm}=0$ (it is the QLC for the constant metric in the lattice in the sense of [2].) Notice that $\Omega^{1}(\mathbb{Z})$ is 2-dimensional over the algebra because $\partial_{\pm}$ are linearly independent as operators. For the same reason the top ‘volume’ form ${\rm Vol}:=e^{+}\wedge e^{-}$ is a 2-form. We fix the above calculus $\Omega(\mathbb{Z})$ and our first task is to extend it to a calculus $\Omega(J^{\infty})$.